OpenStudy (anonymous):

ideal mechanical advantage of 12 input arm length=0.6 m how long is the output arm?

OpenStudy (theeric):

Hi! So, we're working with a lever arm. Right?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

Can anyone help me??

OpenStudy (theeric):

Okay! So, do you know what mechanical advantage means?

OpenStudy (anonymous):

No, just the formula

OpenStudy (anonymous):

When output arm is bigger than input arm?

OpenStudy (theeric):

Okay! Well, even if all you do is algebra, sometimes knowing what things are can help. Mechanical advantage is the ratio \(force\ coming\ from\ the\ machine\) to \(force\ into\ a\ machine\) So, your "MA" is saying that you are getting 12 times as much force out than you put in. So, you do know the formula for a lever?

OpenStudy (anonymous):

For ma?

OpenStudy (anonymous):

What's the difference between IDEAL mechanical advantage and just mechanical advantage

OpenStudy (theeric):

Ideal mechanical advantage is the mechanical advantage you expect if the machine was perfect: no friction, no air resistance, no energy losses due to your metal lever bending, or other issues like that. In a non-ideal case, you have to deal with a non-conservative force. I think that is the only difference, but I'm not completely sure.

OpenStudy (anonymous):

OK, in my question though it as about ideal mechanical advantage

OpenStudy (theeric):

Right, so we don't have to worry about pesky things. We just consider the force out per the force in.

OpenStudy (anonymous):

Ok, so it doesn t matter?

OpenStudy (theeric):

Friction and the other things don't matter - right. So we want \(\dfrac{F_{out}}{F_{in}}\). This out to be \(\dfrac{r_{input\ side}}{r_{output\ side}}\) So, you can either remember that, or be prepared to derive it.

OpenStudy (anonymous):

OK, so how would you solve it?

OpenStudy (theeric):

Substitution! Using the definition of mechanical advantage and then what the ratio of those forces turns out to be: \(MA=\dfrac{F_{out}}{F_{in}}=\dfrac{r_{input\ side}}{r_{output\ side}}\) \(\qquad\Downarrow\) \(12=\dfrac{F_{out}}{F_{in}}=\dfrac{0.6}{\color{red}{r_{output\ side}}}\) Now you can look at it and see how to deal with this algebraically.

OpenStudy (anonymous):

Ohhhh, OK thanks!

OpenStudy (anonymous):

I think I get it now!

OpenStudy (theeric):

Cool! :) Let me know if you want to see \(why\) the ratio of the forces "turns out to be" that. If you have to memorize formulas, I think it's nice to have a way to check if you remembered it correctly.

OpenStudy (anonymous):

OK, thanks for your help!! ;)

OpenStudy (anonymous):

It helped a lot.

OpenStudy (theeric):

You're very welcome! :) I'm glad you asked questions - that always helps.

OpenStudy (anonymous):

Thanks!

OpenStudy (theeric):

:)

OpenStudy (anonymous):

:)